Circular No. 7742

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fd IN UGC 11957 M. Papenkova and W. D. Li, University of California at Berkeley, re-

port the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.4) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Nov. 1.2 and 2.2 UT. SN 2001fd is located at α = 22^{h}14^{m}29^{s}.64, δ = +5^{o}01′37′′.6 (equinox 2000.0), which is 8′′.3 east and 15′′.9 south of the nucleus of UGC 11957. A KAIT image taken on Oct. 26.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001fe IN UGC 5129 G. M. Hurst, Basingstoke, England, reports the discovery of an ap-

parent supernova (mag 16.0) by M. Armstrong, Rolvenden, England on unfiltered CCD images taken on Nov. 2.23 and 3.08 UT in the course of the U.K. Nova/Supernova Patrol, showing the new object at α = 9^{h}37^{m}57^{s}.10, δ = +25^{o}29′41′′.3 (equinox 2000.0), which is 15′′ west and 1′′ south of the center of UGC 5129. An unfiltered CCD image taken by T. Boles, Coddenham, England, on Nov. 3.135 yields position end figures 57^{s}.03, 41′′.7 (and offsets 13′′.5 west, 0′′.1 south). Nothing was present at this location on an image taken by Armstrong on 2000 Nov. 4 (limiting mag 19.0) or on second Palomar Sky Survey images taken on 1995 May 19 (limiting blue mag 22.5) and 1998 Apr. 23 (limiting red mag 20.5).

2001 SL_{9}

P. Pravec and P. Kušnirák, Ondřejov Observatory; and B. Warner, Palmer Divide, CO, write: “Our photometric observations of this Apollotype object (cf. MPEC 2001-S16, MPO 18074) on ten nights during Oct. 11.0–21.3 UT reveal a lightcurve that appears to be a co-addition of two components of different periods: 16.40 ± 0.02 hr (with amplitude 0.10 mag, showing two sharp, narrow minima and little or no variation at phases between them) and 2.4003 ± 0.0003 hr (with amplitude 0.08 mag and most of the signal in the second and fourth harmonics). The lightcurve character indicates that 2001 SL_{9} is a binary system with the long period being a mutual orbital period (the sharp minima being occultation + eclipse events between the binary components), the short period being a rotation period of the primary, and a secondary-to-primary diameter ratio of 0.31 with formal error 0.02 (see Pravec et al. 2000, Icarus 146, 190). Follow-up observations are desirable.”

2001 November 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green